OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of –2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x). Find the polynomial f(x) that has the roots of –2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x). @Mathematics

OpenStudy (anonymous):

help?

hero (hero):

(x-3)(x+2) = x^2-x-6 i don't know what "muliplicity" 2 means

OpenStudy (anonymous):

meaning like there is two of (x-3)

hero (hero):

Okay

OpenStudy (anonymous):

But what do they want in the last question?

hero (hero):

(x-3)(x+2)= 0 x-3 = 0 x+2 = 0 x = 3 x = -2

OpenStudy (anonymous):

ok, but thats verifying it?

hero (hero):

algebraically yes. You can also graph x^2-x-6 to verify as well

OpenStudy (anonymous):

Thank you :)